# MoIIE: Mixture of Intra- and Inter-Modality Experts for Large Vision Language Models

### v1

@@ 3.7 Visualization of Token Pathways
Figure 4: Visualization of experts activated pathways. The figure shows the top-2 activated experts for text and image tokens, with Expert V and Expert L are intra-modality experts, Expert S denotes inter-modality experts.
As illustrated in Figure 4, we visualize the expert activation pathways for image and text tokens during the feed-forward process in the MoIIE module on the MME test set. Our analysis reveals distinct activation patterns across different model layers. In the shallow layers (the first half), both image and text tokens predominantly activate intra-modality experts in their respective modalities (Expert V for vision and Expert L for text). We hypothesize that this behavior is due to shallow layers focusing on extracting modality-specific features with minimal cross-modal interaction.
In contrast, the deeper layers of the model (the second half) show a shift toward activating inter-modality experts (Expert S), promoting deeper cross-modal interactions. Interestingly, both image and text tokens tend to activate the same inter-modality experts in these deep layers. For example, in the 9th MoIIE layer, Expert S_1 is frequently activated, while in the 12th MoIIE layer, Expert S_2 becomes the dominant expert. These findings indicate that our MoIIE module effectively balances modality-specific and cross-modal processing, dynamically adjusting expert activation to facilitate both early-stage feature extraction and late-stage modality fusion.
@@ 4 Related Works
Recent advancements in large language models (LLMs)  Bai et al. (2023a); AI@Meta (2024); Touvron et al. (2023b); Chiang et al. (2023); Jiang et al. (2023); Bi et al. (2024); Team (2023); OpenAI (2023b, a); Touvron et al. (2023a) have demonstrated remarkable generalization and instruction-following capabilities, driving progress toward general-purpose AI. To extend these abilities to both linguistic and visual domains, research on Large Vision-Language Models (LVLMs) has gained increasing attention Li et al. (2023b); Alayrac et al. (2022); Liu et al. (2023a, b, 2024b). Recent efforts in LVLMs have improved performance through high-quality data He et al. (2024); Chen et al. (2024a); Zhang et al. (2024), additional training phases Bai et al. (2023b); Lu et al. (2024a); Laurençon et al. (2024), high-resolution input support Li et al. (2024e); Wang et al. (2024b), and multi-encoder enhancements Shi et al. (2024); Liu et al. (2024a); Fan et al. (2024). The latest open-source LVLMs Li et al. (2024a); Wang et al. (2024a); Chen et al. (2025) leverage dynamic resolution, large-scale datasets, and advanced LLMs to achieve state-of-the-art results.
@@
The MoE paradigm Jacobs et al. (1991) enables selective expert activation, allowing for efficient scaling while maintaining manageable computational costs. Integrating MoE into LLMs involves replacing feed-forward networks with expert modules, using a Top-K router for activation Lepikhin et al. (2020); Du et al. (2022); FedusF et al. (2022); Zoph et al. (2022); Rajbhandari et al. (2022); Xue et al. (2024). Sparse upcycling Komatsuzaki et al. (2023) reduces costs by transforming dense models into sparse ones. Recent MoE models xAI (2024); MistralAITeam (2023); Muennighoff et al. (2024); DeepSeek-AI (2024a, b) refine performance through large-scale data and enhanced training strategies.
@@
MoE is used in LVLMs, such as ARIA Li et al. (2025) and DeepSeek-VL2 Wu et al. (2024), leveraging pre-trained MoE-based LLM backbones MistralAITeam (2023); DeepSeek-AI (2024a) but are limited by fixed expert settings and routing mechanisms along with high training costs. Sparse upcycling as in  Chen et al. (2024b); Lin et al. (2024a, c) are utilized to improve scalability but still faces challenges in capturing both modality-specific features and cross-modal relationships. Zhou et al. (2025) introduce extend the traditional Parameter-efficient fine-tuning framework to support multi-modal expert, but the impact of the modality expert groups has not been thoroughly explored. To address these issues, we propose a modality-aware MoE architecture, which models both intra-modality experts and inter-modal experts. Our two-stage training strategy optimizes MoE activation, offering a more flexible and scalable approach for building high-performance LVLMs.
@@ 5 Conclusion
In this study, we propose MoIIE, a novel Mixture-of-Experts architecture paired with a simple yet effective two-stage training strategy that enables the flexible construction of powerful MoE-based LVLMs from any dense LLM. MoIIE organizes experts into specialized intra-modality and shared inter-modality groups to create dedicated pathways for both visual and linguistic processing while enabling rich cross-modal associations, significantly enhancing performance across a wide range of multi-modal tasks. Extensive experiments demonstrate that our framework achieves superior scaling efficiency compared to existing dense and sparse LVLM architectures.
Broader Impacts and Limitations
@@ Appendix A Appendix
Learning rate
1e-3
{connection module}
2e-5
{LLM,connection module}
2e-6
{visual encoder}

### v2

@@ 3.7 Visualization of Token Pathways
Figure 4: Visualization of experts activated pathways. The figure shows the top-2 activated experts for text and image tokens, with Expert V and Expert L are intra-modality experts, Expert S are inter-modality experts.
As shown in Figure 4, we visualize expert activation pathways for image and text tokens in the MoIIE module on the MME test set. Distinct layer-wise patterns emerge: in shallow layers, tokens primarily activate intra-modality experts (Expert V for vision, Expert L for text), reflecting a focus on modality-specific feature extraction with limited cross-modal interaction. In deeper layers, activation shifts toward inter-modality experts (Expert S), enabling stronger cross-modal fusion. Notably, both modalities often converge on the same inter-modality experts (e.g., Expert S_1 in the 9th layer, Expert S_2 in the 12th). These results demonstrate that MoIIE dynamically balances modality-specific and cross-modal processing, supporting early-stage representation learning and late-stage integration.
@@ 4 Related Works
Recent advances in LLMs Bai et al. (2023a); AI@Meta (2024); Touvron et al. (2023b); Chiang et al. (2023); Jiang et al. (2023); Bi et al. (2024); Team (2023); OpenAI (2023b, a); Touvron et al. (2023a) have shown strong generalization and instruction-following abilities, spurring research on Large Vision-Language Models (LVLMs) Li et al. (2023b); Alayrac et al. (2022); Liu et al. (2023a, b, 2024b). Progress in LVLMs has been driven by high-quality data He et al. (2024); Chen et al. (2024a); Zhang et al. (2024), extended training schemes Bai et al. (2023b); Lu et al. (2024a); Laurençon et al. (2024), support for high-resolution inputs Li et al. (2024e); Wang et al. (2024b), and multi-encoder designs Shi et al. (2024); Liu et al. (2024a); Fan et al. (2024). The latest open-source LVLMs Li et al. (2024a); Wang et al. (2024a); Chen et al. (2025) achieve state-of-the-art results by integrating large-scale datasets, dynamic resolution, and advanced LLMs.
@@
The MoE paradigm Jacobs et al. (1991) scales model capacity by activating only a subset of experts, balancing efficiency and performance. It typically replaces feed-forward layers with expert modules and relies on Top-K routing Lepikhin et al. (2020); Du et al. (2022); FedusF et al. (2022); Zoph et al. (2022); Rajbhandari et al. (2022); Xue et al. (2024). Sparse upcycling Komatsuzaki et al. (2023) further reduces costs by converting dense models into sparse ones. Recent MoE-based LLMs xAI (2024); MistralAITeam (2023); Muennighoff et al. (2024); DeepSeek-AI (2024a, b) improve stability and efficiency through large-scale training strategies.
@@
MoE has also been applied to LVLMs, e.g., ARIA Li et al. (2025) and DeepSeek-VL2 Wu et al. (2024), which leverage MoE-based LLM backbones MistralAITeam (2023); DeepSeek-AI (2024a) but remain constrained by fixed expert settings and high training costs. Sparse upcycling Chen et al. (2024b); Lin et al. (2024a, c) improves scalability, yet struggles to capture both modality-specific and cross-modal interactions. More recent work Zhou et al. (2025) extends parameter-efficient fine-tuning to multimodal experts, but the role of modality-specific expert groups remains underexplored. To address these gaps, we propose a modality-aware MoE architecture that integrates intra- and inter-modality experts. Combined with a two-stage training strategy, our approach enables flexible expert allocation and more scalable, high-performance LVLMs.
@@ 5 Conclusion
In this study, we propose MoIIE, a novel Mixture-of-Experts architecture paired with a simple yet effective two-stage training strategy that enables the flexible construction of powerful MoE-based LVLMs from any dense LLM. MoIIE organizes experts into specialized intra-modality and shared inter-modality groups to create dedicated pathways for both visual and linguistic processing while enabling rich cross-modal associations, significantly enhancing performance across a wide range of multi-modal tasks. Extensive experiments demonstrate that our framework achieves superior scaling efficiency and performance compared to existing dense and sparse LVLM architectures.
Limitations
@@ A.2 Training Details
Learning rate 1e-3 {connection module}
2e-5 {LLM,connection module}
2e-6 {visual encoder}
@@ A.3 Expert Analysis
[…]
Data Visual Encoder General Multi-modal QA Hallucination OCR-based QA Knowledge-based QA AVG
[…]
Dense 2.7M CLIP-L-336 74.5 63.0 81.3 40.9 69.9 86.5 31.2 64.1 47.7 57.1 73.2 39.6 30.6 58.4
Vanilla MoE 2.7M CLIP-L-336 74.6 63.1 81.1 41.0 69.7 86.5 30.9 64.3 47.5 57.0 74.0 40.5 31.0 58.5
Modality MoE 2.7M CLIP-L-336 74.0 63.0 81.5 39.7 69.5 86.6 32.3 64.4 48.1 57.7 73.7 40.1 31.3 58.6
[…]
Dense 4M SigLIP-So400m-384 76.2 64.6 82.5 42.3 71.4 88.1 32.4 66.1 49.4 58.8 75.0 41.3 32.1 60.0
Vanilla MoE 4M SigLIP-So400m-384 76.4 64.9 82.9 42.9 71.6 88.9 33.8 66.2 50.2 59.7 75.8 43.0 33.3 60.7
Modality MoE 4M SigLIP-So400m-384 75.1 64.7 82.7 41.9 71.3 88.0 35.3 67.4 50.8 59.0 75.3 41.8 32.5 60.4
[…]
A.4 More Experimental Results
we conduct extended ablation experiments using a different vision encoder(CLIP-L-336) as shown in Table A.3, replacing the vision encoder does not affect our main conclusions. Under identical training conditions,the proposed MoIIE architecture consistently outperforms all other ablation variants across most of the 13 benchmarks and shows a clear advantage in terms of average performance.
We also adopt a larger dataset that matches the scale of modern open multimodal benchmarks. Specifically, we followed LLaVA-OV Li et al. (2024a), which provides the high overall data quality among these open datasets a total training corpus of approximately 8M samples, 4M for insturct tuing stage, and used it as our expanded training corpus.
As shown in Table A.3, using a higher-quality and larger-scale dataset but still identical training conditions across all variants, our proposed MoIIE architecture continues to outperform all other variants across all the 13 multimodal benchmarks. Moreover, MoIIE achieves a more pronounced improvement in average performance (+1.2%) relative to the other variants. with the trainig data more advanced, the performance gap between MoIIE and other architectures widening.
These results demonstrate that MoIIE not only maintains its advantages under stronger training condition, but also scales more effectively in multimodal training, confirming its robustness and generality.